OpenStudy (anonymous):

What value is added to both sides of the equation x2 - 8x = -10 in order to solve by completing the square?

OpenStudy (anonymous):

We add : \[\left(\frac82\right)^2\] In general we complete : \[x^2+ax=b\] to the square by adding : \[\left(\frac a2\right)^2\] to the both sides, so we get : \[x^2+ax+\left(\frac a2\right)^2=b+\left(\frac a2\right)^2\] So : \[\left(x+\frac a2\right)^2=b+\left(\frac a2\right)^2\]

OpenStudy (anonymous):

So is the answer 4? @Noura11

OpenStudy (anonymous):

Actually 16?

OpenStudy (anonymous):

Yes, it is 16 ;)

OpenStudy (anonymous):

Thank you much, I'm going to give you a mental for helping me. haha